# cuTeSpMM: Accelerating Sparse-Dense Matrix Multiplication using GPU Tensor Cores

### v1

@@ 1. Introduction
$Emilia\_923$
$NotreDame\_www$
@@
Can these high throughput, fixed size, dense tensor core hardware implementations be used for sparse-dense matrix multiplication (SpMM)? While sparse matrices used in scientific computing and data analytics tend to be extremely sparse (with far less than 1% of elements being non-zero), the distribution of non-zeros is generally not uniformly random across the 2D index space. The non-zero distribution pattern tends to exhibit clustering in sub-regions of the 2D index space, which leaves open the possibility of developing zero filling strategies to leverage these dense TC Units. Fig. 1 shows the non-zero pattern for two sparse matrices from the SuiteSparse matrix collection (davis2011university, ; suitesparse, ). The matrix $Emilia\_923$ arose from a structural application and has $923136$ rows/columns and $41005206$ nonzero elements (i.e., it is extremely sparse with a density of less than $0.005\%$). However the non-zero elements tend to be clustered around the diagonal. If the 2D $923136\times 923136$ index space is partitioned into a regular grid of $16\times 4$ “bricks” (this is one of the supported dense matrix sizes in Nvidia’s WMMA API for using its tensor cores for matrix-matrix multiplication), the average number of non-zero elements in non-empty bricks is nearly 20% dense. $NotreDame\_www$ is a web graph with $325760$ rows/columns and $929849$ nonzero elements, with an extremely low density of less than $0.0009\%$. Its non-zeros are scattered over a much wider range of the 2D index space, but even for this much sparser and more scattered matrix, the average density of non-zero elements in non-empty $16\times 4$ bricks is nearly 10%. Therefore, there is the potential to develop algorithms to leverage dense tensor core hardware to accelerate sparse matrix computations.
@@ 2. Background
Tensor core matrix multiply-accumulate (MMA) operations are warp-synchronous, meaning all threads within a warp (i.e., a group of 32 threads) collectively participate in the operation. Each thread is responsible for computing a portion of the output matrix $D$ using rectangular slices of matrices $A$ and $B$ (andrewkerr, ). The supported sizes of the matrix subsets are architecture dependent. For example, the Ampere $TF32$ tensor core MMA supports an $A$ shape of ${(16,4)}$, a $B$ shape of ${(4,8)}$, and an output matrix ($D$) shape of ${(16,8)}$. During tensor core MMA operations, every consecutive groups of 4 threads form a thread group within a warp, resulting in 8 thread groups. Each thread within a group loads an element from a row of the A operand matrix. Given the A operand matrix’s 16 rows, each thread group cyclically loads two rows. For the B matrix operand, each thread group loads a column from the B matrix into its registers. Regarding the C operand matrix, each thread group stores two rows of the resulting matrix. Within each row of C, every thread in a group holds two elements.
Developers can use tensor cores via the high level $cublas.gemmex$ library API or by using the lower level $nvcuda::wmma::*$ namespace API from the $mma.h$ cuda header file. To implement a basic Matrix Multiply one can divide the output between the warps such that each warp is responsible for generating a small tile of the output matrix. For each output tile the relevant input tiles are distributed amongst the threads of the warp in which each thread is owning a fraction.
@@ 3.1. Notations
In this paper, we use the following notation: $A$ denotes the sparse input matrix, $B$ denotes the dense input matrix, and $C$ denotes the dense output matrix. $M$ and $K$ represent the number of rows and columns of matrix $A$, respectively. $N$ represents the width of matrix $B$. We refer to a contiguous set of rows of $A$ as a row-panel, and we denote the size of each row panel as $TM$. Each row panel is partitioned along the column dimension into blocks of shape $(TM,TK)$, where $TK$ is the tile size along the column dimension. Each block is further sub-divided into bricks of shape $(brick\_m,brick\_k)$. The matrix $B$ is divided into bricks of shape $(brick\_k,brick\_n)$. Each tensor core operation computes the product of $A[brick\_m][brick\_k]$ with $B[brick\_k][brick\_n]$ and accumulates the results to the output brick $C[brick\_m][brick\_n]$.
@@ 3.2. HRPB Sparse Matrix Data structure
The first step in creating the HRPB is to collect (re-order) all active columns in a row panel and place them together towards the left. A column in a row panel is considered active if it contains at least one non-zero element. These columns are then grouped to form bricks. This process is illustrated in the “compacting“ step of Figure 3 (a). Note that this step reduces the number of active blocks and thus the total operations. Each row panel is divided into blocks of shape $(TM,TK)$, where we set $TM$ either 16 or 32. The $TK$ is set to 16 based on our experiments. On Nvidia Ampere architecture, for TF32 data type, the tensor core operations support fragments (slices) of shape (16, 4) for the A matrix, shape (4, 8) for the B matrix, and the output shape is (16, 8) of the C matrix. Hence, we further subdivide the blocks into bricks of shape ${(brick_{m},brick_{k})}$, where $brick_{m}$ is set to 16, $brick_{k}$ is set to 4, and $brick_{n}$ is set to 8. The compaction step helps to reduce the number of active blocks and bricks, which in turn minimizes the number of FLOPS. To maintain the column index information after compacting, we maintain an array called ${ActiveColIdx[]}$ for each row panel. The $i$-th element of this array represents the original column id of the $i$-th active column.
@@
The active bricks in a block are stored in a column-major format, which closely resembles the Compressed Sparse Column (CSC) format. The “To BlkCSC” component of Figure 3 (a) illustrates our data structure. For each active brick, we encode the non-zero distribution pattern, i.e., the indices corresponding to non-zero elements, into a 64-bit vector. Since the brick size is 64, every bit in the 64-bit vector indicates the corresponding position of the brick is non-zero or not. For instance, pattern[i] indicates if the ith element in the brick is non-zero or not. Figure 3 (b) illustrates this process. These patterns are stored in an array called ‘patterns’. Additionally, we utilize an array named ${nnz\_array}$ to maintain the non-zero values across different active bricks in CSC order. The non-zero values within each brick are ordered in row major and appended to the ${nnz\_array}$ of block. In addition to the ${nnz\_array}$ and ${patterns}$, we utilize a ’rows’ array to indicate the row ID of each active brick and a ’colPtr’ array to denote the starting index of each brick column. Figure 4 illustrates the structure of a block. The ’colPtr’, ’patterns’, and ’rows’ arrays are collectively referred to as metadata.
@@
Figure 5 illustrates the HRPB data structure, which organizes all the blocks in the sparse matrix. The ‘packedBlocks’ in the HRPB data structure represent a memory chunk containing all the blocks of all the row panels packed together. Different row panels may contain different number of blocks. To identify all blocks belonging to individual row panels, we utilize an array called ‘blockedRowPtr’. This array has a size of ${\frac{M}{TM}+1}$ and resembles the row pointer array in the CSR format, storing the starting block index of each row panel. For instance, blockedRowPtr[i] indicates the starting block index of row panel i, and blockedRowPtr[i+1] - blockedRowPtr[i] indicates the number of blocks inside row panel ‘i’. However, since each block may contain a varying number of non-zero elements, their memory requirement may differ. Hence, ‘blockedRowPtr’ alone cannot precisely locate the position of a block within the ‘packedBlocks’ memory chunk. Hence, we use sizePtr$[NUM\_BLKS+1]$ to determine the starting memory address of each block in ‘packedBlocks’. The ‘activeCols’ array stores the active column indices of each row panel.
1 Input: blockedRowPtr[], sizePtr[], packedBlocks[], numRowPanels
@@ 2 Output: matrix C[]
3 shared ${SM\_A[TM*TK],SM\_B[TK][N]}$
4 ${a\_frag}[2]$
5 ${b\_frag}[4]$
@@ 6 ${c\_frag}[4][TM/brick_{m}][4]$
7 ${warp\_id=threadIdx/32}$
@@ 8 ${lane\_id=threadIdx\%32}$
9 ${group\_id=lane\_id/4}$
@@ 10 ${tid\_in\_group=lane\_id\%4}$
11 ${row\_panel=threadblock\_id}$
12 ${block\_start=blockedRowPtr[row\_panel]}$
13 ${block\_end=blockedRowPtr[row\_panel+1]}$
14 for $block\_id=block\_start$ to $block\_end$ do
15       $block\_addr=sizePtr[block\_id]$
16       $block\_addr\_end=sizePtr[block\_id+1]$
17       $SM\_A=packedBlocks[block\_addr:block\_addr\_end]$
18       $block=(Block)SM\_A$
19       for $i=warp\_id$ to TK step $num\_warps$ do
20             $b\_row=block.active\_cols[i]$
21             for $j=lane\_id$ to N step 32 do
22                   $SM\_B[i][j]=B[b\_row*N+j]$
@@ 28                   $b\_frag[n]=SM\_B[i*brick_{k}+tid\_group]$
29                   $[warp\_id*4*brick_{n}+n*brick_{n}+group\_id]$
@@ 31                   $afrag[0]=afrag[1]=0$
40                  $nnz\_offset+=count\_1s(pattern)$
@@
42                         $warp\_wmma(a\_frag,b\_frag[n],$
43                         $c\_frag[n][row])$
@@ 50 Write(C[m: m+ TM][n:n+$brick_{n}*4$],$c\_frag$)
For small feature sizes, like 32, allocating a thread block of size 128 to process each block would lead to inefficient resource utilization. Therefore, for smaller feature sizes, we decrease the number of warps in a thread block to ${\frac{N}{warp\_coarsening\_level\times 8}}$ (where the 8 represents the tensor core MMA operand tile size along N, referred to as $brick\_n$). For large N size, a row panel is assigned multiple thread blocks, while each thread block produce a ${TM\times 128}$ portion of the output matrix. For simplicity, for the later discussion, we will assume that N is 128, which means the thread block size is 128.
@@
Algorithm 1 illustrates the kernel design. Each thread sequentially iterates over the blocks along the reduction dimension K inside the row panel (Line 14). For each block, we initially load the non-zeros of the active bricks inside the block and the corresponding meta-data into shared memory $SM_{A}$ in a coalesced manner with the assistance of sizePtr (line 17). Next, we unpack A in line 18. Given that the maximum non-zero inside the block is ${TM\times TK}$, within the GPU thread block, we allocate ${TM\times TK+MetaDataSize}$ shared memory space to store the block. Following that, we load the required rows of B into shared memory. We utilize the ${block.active\_cols}$ array with a size of TK to identify the necessary rows of B. We fetch the corresponding rows from global memory in a coalesced manner and pack them into a contiguous 2D buffer in shared memory (Lines 19-22). Fig. 6 illustrates the process of loading the necessary rows of B into the shared memory buffer.
@@
After loading the necessary rows of B to shared memory, we will traverse the active bricks inside the block in a CSC manner. Inside a block, there will be ${\frac{TK}{brick_{k}}}$brick columns. For each brick column, the block.colPtr indicates the starting and ending active brick index of this brick column. The block.rows pointer is used to figure out the brick row index and the brick.patterns is to find out the brick non-zero pattern by a given active brick index. As we explained earlier, the non-zero brick pattern is a 64-bit vector. In line 34, each thread in a warp calculates the number of non-zero elements assigned to the threads with a lane id less than the current thread. This index represents the position of the non-zero element that this thread should load in the current brick. Within the block, multiple bricks will exist. To keep track of the starting position of non-zeros in each active brick, we introduce the ${nnz\_offset}$ variable. Initially set to 0, ${nnz\_offset}$ is incremented after decoding a brick pattern by the number of non-zeros in the brick. Thus, ${nnz\_offset}$ indicates the starting position of non-zeros for the next brick to be processed. The scanning of prefix bits is highly efficient and is implemented using bitwise operations and is computed by the CUDA cores, which do not participate in the core MMA computations (Tensor Cores handle these computations).
@@
After reading the corresponding non-zeros into registers, a warp collectively processes the four distinct output tiles of the C matrix within the loop (line 29). Inside the loop, each thread in the warp is required to load the corresponding element from the B matrix in shared memory, denoted as $SM_{B}$, into registers. Following this, the thread performs the MMA operation and stores the partial results in ${cfrag}$. Once all the blocks in a row panel have been iterated, each thread will write its ${cfrag}$ to global memory. This implies that the thread block will collectively write out a ${TM\times 128}$ tile of the C matrix to global memory.

### v2

@@ 1. Introduction
Emilia_923Emilia\_923
NotreDame_wwwNotreDame\_www
@@
Can these high throughput, fixed size, dense tensor core hardware implementations be used for sparse-dense matrix multiplication (SpMM)? While sparse matrices used in scientific computing and data analytics tend to be extremely sparse (with far less than 1% of elements being non-zero), the distribution of non-zeros is generally not uniformly random across the 2D index space. The non-zero distribution pattern tends to exhibit clustering in sub-regions of the 2D index space, which leaves open the possibility of developing zero filling strategies to leverage these dense TC Units. Fig. 1 shows the non-zero pattern for two sparse matrices from the SuiteSparse matrix collection (davis2011university, ; suitesparse, ). The matrix Emilia_923Emilia\_923 arose from a structural application and has $923136$ rows/columns and $41005206$ nonzero elements (i.e., it is extremely sparse with a density of less than $0.005\%$). However the non-zero elements tend to be clustered around the diagonal. If the 2D $923136\times 923136$ index space is partitioned into a regular grid of $16\times 4$ “bricks” (this is one of the supported dense matrix sizes in Nvidia’s WMMA API for using its tensor cores for matrix-matrix multiplication), the average number of non-zero elements in non-empty bricks is nearly 20% dense. NotreDame_wwwNotreDame\_www is a web graph with $325760$ rows/columns and $929849$ nonzero elements, with an extremely low density of less than $0.0009\%$. Its non-zeros are scattered over a much wider range of the 2D index space, but even for this much sparser and more scattered matrix, the average density of non-zero elements in non-empty $16\times 4$ bricks is nearly 10%. Therefore, there is the potential to develop algorithms to leverage dense tensor core hardware to accelerate sparse matrix computations.
@@ 2. Background
Tensor core matrix multiply-accumulate (MMA) operations are warp-synchronous, meaning all threads within a warp (i.e., a group of 32 threads) collectively participate in the operation. Each thread is responsible for computing a portion of the output matrix $D$ using rectangular slices of matrices $A$ and $B$ (andrewkerr, ). The supported sizes of the matrix subsets are architecture dependent. For example, the Ampere $TF32$ tensor core MMA supports an $A$ shape of (16,4){(16,4)}, a $B$ shape of (4,8){(4,8)}, and an output matrix ($D$) shape of (16,8){(16,8)}. During tensor core MMA operations, every consecutive groups of 4 threads form a thread group within a warp, resulting in 8 thread groups. Each thread within a group loads an element from a row of the A operand matrix. Given the A operand matrix’s 16 rows, each thread group cyclically loads two rows. For the B matrix operand, each thread group loads a column from the B matrix into its registers. Regarding the C operand matrix, each thread group stores two rows of the resulting matrix. Within each row of C, every thread in a group holds two elements.
Developers can use tensor cores via the high level $cublas.gemmex$ library API or by using the lower level nvcuda::wmma::nvcuda::wmma::* namespace API from the $mma.h$ cuda header file. To implement a basic Matrix Multiply one can divide the output between the warps such that each warp is responsible for generating a small tile of the output matrix. For each output tile the relevant input tiles are distributed amongst the threads of the warp in which each thread is owning a fraction.
@@ 3.1. Notations
In this paper, we use the following notation: $A$ denotes the sparse input matrix, $B$ denotes the dense input matrix, and $C$ denotes the dense output matrix. $M$ and $K$ represent the number of rows and columns of matrix $A$, respectively. $N$ represents the width of matrix $B$. We refer to a contiguous set of rows of $A$ as a row-panel, and we denote the size of each row panel as $TM$. Each row panel is partitioned along the column dimension into blocks of shape $(TM,TK)$, where $TK$ is the tile size along the column dimension. Each block is further sub-divided into bricks of shape (brick_m,brick_k)(brick\_m,brick\_k). The matrix $B$ is divided into bricks of shape (brick_k,brick_n)(brick\_k,brick\_n). Each tensor core operation computes the product of A[brick_m][brick_k]A[brick\_m][brick\_k] with B[brick_k][brick_n]B[brick\_k][brick\_n] and accumulates the results to the output brick C[brick_m][brick_n]C[brick\_m][brick\_n].
@@ 3.2. HRPB Sparse Matrix Data structure
The first step in creating the HRPB is to collect (re-order) all active columns in a row panel and place them together towards the left. A column in a row panel is considered active if it contains at least one non-zero element. These columns are then grouped to form bricks. This process is illustrated in the “compacting“ step of Figure 3 (a). Note that this step reduces the number of active blocks and thus the total operations. Each row panel is divided into blocks of shape $(TM,TK)$, where we set $TM$ either 16 or 32. The $TK$ is set to 16 based on our experiments. On Nvidia Ampere architecture, for TF32 data type, the tensor core operations support fragments (slices) of shape (16, 4) for the A matrix, shape (4, 8) for the B matrix, and the output shape is (16, 8) of the C matrix. Hence, we further subdivide the blocks into bricks of shape $(brick_{m},brick_{k})$}, where $brick_{m}$ is set to 16, $brick_{k}$ is set to 4, and $brick_{n}$ is set to 8. The compaction step helps to reduce the number of active blocks and bricks, which in turn minimizes the number of FLOPS. To maintain the column index information after compacting, we maintain an array called ActiveColIdx[]{ActiveColIdx[]} for each row panel. The $i$-th element of this array represents the original column id of the $i$-th active column.
@@
The active bricks in a block are stored in a column-major format, which closely resembles the Compressed Sparse Column (CSC) format. The “To BlkCSC” component of Figure 3 (a) illustrates our data structure. For each active brick, we encode the non-zero distribution pattern, i.e., the indices corresponding to non-zero elements, into a 64-bit vector. Since the brick size is 64, every bit in the 64-bit vector indicates the corresponding position of the brick is non-zero or not. For instance, pattern[i] indicates if the ith element in the brick is non-zero or not. Figure 3 (b) illustrates this process. These patterns are stored in an array called ‘patterns’. Additionally, we utilize an array named nnz_array{nnz\_array} to maintain the non-zero values across different active bricks in CSC order. The non-zero values within each brick are ordered in row major and appended to the nnz_array{nnz\_array} of block. In addition to the nnz_array{nnz\_array} and patterns{patterns}, we utilize a ’rows’ array to indicate the row ID of each active brick and a ’colPtr’ array to denote the starting index of each brick column. Figure 4 illustrates the structure of a block. The ’colPtr’, ’patterns’, and ’rows’ arrays are collectively referred to as metadata.
@@
Figure 5 illustrates the HRPB data structure, which organizes all the blocks in the sparse matrix. The ‘packedBlocks’ in the HRPB data structure represent a memory chunk containing all the blocks of all the row panels packed together. Different row panels may contain different number of blocks. To identify all blocks belonging to individual row panels, we utilize an array called ‘blockedRowPtr’. This array has a size of $\frac{M}{TM}+1$} and resembles the row pointer array in the CSR format, storing the starting block index of each row panel. For instance, blockedRowPtr[i] indicates the starting block index of row panel i, and blockedRowPtr[i+1] - blockedRowPtr[i] indicates the number of blocks inside row panel ‘i’. However, since each block may contain a varying number of non-zero elements, their memory requirement may differ. Hence, ‘blockedRowPtr’ alone cannot precisely locate the position of a block within the ‘packedBlocks’ memory chunk. Hence, we use sizePtr[NUM_BLKS+1][NUM\_BLKS+1] to determine the starting memory address of each block in ‘packedBlocks’. The ‘activeCols’ array stores the active column indices of each row panel.
1Input: blockedRowPtr[], sizePtr[], packedBlocks[], numRowPanels
@@ 2 Output: matrix C[]
3 shared SM_A[TMTK],SM_B[TK][N]{SM\_A[TM*TK],SM\_B[TK][N]}
4 a_frag[2]{a\_frag}[2]
5 b_frag[4]{b\_frag}[4]
@@ 6 ${c\_frag}[4][TM/brick_{m}][4]$
7 warp_id=threadIdx/32{warp\_id=threadIdx/32}
@@ 8 $lane\_id=threadIdx\%32$}
9 group_id=lane_id/4{group\_id=lane\_id/4}
@@ 10 $tid\_in\_group=lane\_id\%4$}
11 row_panel=threadblock_id{row\_panel=threadblock\_id}
12 block_start=blockedRowPtr[row_panel]{block\_start=blockedRowPtr[row\_panel]}
13 block_end=blockedRowPtr[row_panel+1]{block\_end=blockedRowPtr[row\_panel+1]}
14 for block_id=block_startblock\_id=block\_start to block_endblock\_end do
15  block_addr=sizePtr[block_id]block\_addr=sizePtr[block\_id]
16  block_addr_end=sizePtr[block_id+1]block\_addr\_end=sizePtr[block\_id+1]
17  SM_A=packedBlocks[block_addr:block_addr_end]SM\_A=packedBlocks[block\_addr:block\_addr\_end]
18  block=(Block)SM_Ablock=(Block)SM\_A
19  for i=warp_idi=warp\_id to TK step num_warpsnum\_warps do
20     b_row=block.active_cols[i]b\_row=block.active\_cols[i]
21     for j=lane_idj=lane\_id to N step 32 do
22        SM_B[i][j]=B[b_rowN+j]SM\_B[i][j]=B[b\_row*N+j]
@@ 28        $b\_frag[n]=SM\_B[i*brick_{k}+tid\_group]$
29        $warp\_id*4*brick_{n}+n*brick_{n}+group\_id]$
@@ 31        $afrag[0]=afrag[1]=0$
40       nnz_offset+=count_1s(pattern)nnz\_offset+=count\_1s(pattern)
@@
42           warp_wmma(a_frag,b_frag[n],warp\_wmma(a\_frag,b\_frag[n],
43           c_frag[n][row])c\_frag[n][row])
@@
For small feature sizes, like 32, allocating a thread block of size 128 to process each block would lead to inefficient resource utilization. Therefore, for smaller feature sizes, we decrease the number of warps in a thread block to $\frac{N}{warp\_coarsening\_level\times 8}$} (where the 8 represents the tensor core MMA operand tile size along N, referred to as brick_nbrick\_n). For large N size, a row panel is assigned multiple thread blocks, while each thread block produce a $TM\times 128$} portion of the output matrix. For simplicity, for the later discussion, we will assume that N is 128, which means the thread block size is 128.
@@
Algorithm 1 illustrates the kernel design. Each thread sequentially iterates over the blocks along the reduction dimension K inside the row panel (Line 14). For each block, we initially load the non-zeros of the active bricks inside the block and the corresponding meta-data into shared memory $SM_{A}$ in a coalesced manner with the assistance of sizePtr (line 17). Next, we unpack A in line 18. Given that the maximum non-zero inside the block is $TM\times TK$}, within the GPU thread block, we allocate $TM\times TK+MetaDataSize$} shared memory space to store the block. Following that, we load the required rows of B into shared memory. We utilize the block.active_cols{block.active\_cols} array with a size of TK to identify the necessary rows of B. We fetch the corresponding rows from global memory in a coalesced manner and pack them into a contiguous 2D buffer in shared memory (Lines 19-22). Fig. 6 illustrates the process of loading the necessary rows of B into the shared memory buffer.
@@
After loading the necessary rows of B to shared memory, we will traverse the active bricks inside the block in a CSC manner. Inside a block, there will be $\frac{TK}{brick_{k}}$}brick columns. For each brick column, the block.colPtr indicates the starting and ending active brick index of this brick column. The block.rows pointer is used to figure out the brick row index and the brick.patterns is to find out the brick non-zero pattern by a given active brick index. As we explained earlier, the non-zero brick pattern is a 64-bit vector. In line 34, each thread in a warp calculates the number of non-zero elements assigned to the threads with a lane id less than the current thread. This index represents the position of the non-zero element that this thread should load in the current brick. Within the block, multiple bricks will exist. To keep track of the starting position of non-zeros in each active brick, we introduce the nnz_offset{nnz\_offset} variable. Initially set to 0, nnz_offset{nnz\_offset} is incremented after decoding a brick pattern by the number of non-zeros in the brick. Thus, nnz_offset{nnz\_offset} indicates the starting position of non-zeros for the next brick to be processed. The scanning of prefix bits is highly efficient and is implemented using bitwise operations and is computed by the CUDA cores, which do not participate in the core MMA computations (Tensor Cores handle these computations).
@@
After reading the corresponding non-zeros into registers, a warp collectively processes the four distinct output tiles of the C matrix within the loop (line 29). Inside the loop, each thread in the warp is required to load the corresponding element from the B matrix in shared memory, denoted as $SM_{B}$, into registers. Following this, the thread performs the MMA operation and stores the partial results in cfrag{cfrag}. Once all the blocks in a row panel have been iterated, each thread will write its cfrag{cfrag} to global memory. This implies that the thread block will collectively write out a $TM\times 128$} tile of the C matrix to global memory.